### v1

# Domain-Specific Code Language Models: Unraveling the Potential for HPC Codes and Tasks

Abstract
With easier access to powerful compute resources, there is a growing trend in AI for software development to develop larger language models (LLMs) to address a variety of programming tasks. Even LLMs applied to tasks from the high-performance computing (HPC) domain are huge in size and demand expensive compute resources for training. This is partly because these LLMs for HPC tasks are obtained by finetuning existing LLMs that support several natural and/or programming languages. We found this design choice confusing — why do we need large LMs trained on natural languages and programming languages unrelated to HPC for HPC-specific tasks?
In this line of work, we aim to question choices made by existing LLMs by developing smaller LMs for specific domains — we call them domain-specific LMs. Specifically, we start with HPC as a domain and build an HPC-specific LM, named MonoCoder, which is orders of magnitude smaller than existing LMs but delivers similar, if not better performance, on non-HPC and HPC tasks. Specifically, we pre-trained MonoCoder on an HPC-specific dataset (named HPCorpus) of C and C++ programs mined from GitHub. We evaluated the performance of MonoCoder against conventional multi-lingual LLMs. Results demonstrate that MonoCoder, although much smaller than existing LMs, achieves similar results on normalized-perplexity tests and much better ones in CodeBLEU competence for high-performance and parallel code generations. Furthermore, fine-tuning the base model for the specific task of parallel code generation (OpenMP parallel for pragmas) demonstrates outstanding results compared to GPT, especially when local misleading semantics are removed by our novel pre-processor Tokompiler, showcasing the ability of domain-specific models to assist in HPC-relevant tasks.
The sources of this work are available at our GitHub MonoCoder repository.
I Introduction
Recent breakthroughs in the field of AI have led significant attention to language models (LMs) due to their remarkable capabilities in natural language processing [4] and understanding. Large language models (LLMs) [5], particularly exemplified by models such as GPT-3 [6] and its successors [7], have demonstrated the potential to grasp intricate linguistic structure and context, sparking exploration of their applicability beyond natural language tasks. Recently, GPT-3.5 and GPT-4 have been deployed in conversational environments in the form of ChatGPT [8].
In parallel, the field of high-performance computing (HPC) has been tackling increasingly complex and data-intensive problems [9]. The field of HPC has experienced notable advancements in hardware, software, and algorithms, resulting in substantial improvements in computational performance and efficiency [10, 11]. Combining the two trends, integrating LMs into HPC workflows has emerged as a compelling avenue for innovation [12]. For instance, several recent efforts have applied LLMs for automatically inserting OpenMP pragmas or MPI functions in code [13, 14, 15, 16, 17, 18], overcoming limitations of static tools [19, 20, 21, 22, 23].
Although existing LLMs have shown remarkable results on programming tasks [24, 25], such as code generation, bug fixing, etc., to our surprise, we found that they performed poorly on some of HPC-related programming tasks such as code parallelization, vectorization, etc. [26, 15, 16, 12]. This led us to ponder on several research questions: (i) How well do existing LLMs perform on domain-specific tasks such as those for HPC domain, (ii) Will more domain-specific training data help in improving the performance?, (iii) When applying LMs for a specific domain, do we finetune existing LMs, or should we train them from scratch? This last question is specifically important in the context of training costs. For instance, HPC-Coder [16], a recently-introduced LM for HPC tasks, is obtained by finetuning PolyCoder [27] on an HPC dataset; PolyCoder itself is a code LM (not specific to HPC) that is trained on 249 GB of code corpus that is made up of programs written in 12 programming languages. HPC-GPT [28], LM4HPC [12], and AutoParLLM [29] rely on LLaMA as their base model — with up to 34 billion parameters — or even on GPT-4. Such setups looked counter-intuitive to us — Is it not enough to train an LM on HPC-specific languages only? In other words, why do we need an LM trained on Java or Python — with tens or hundreds of billion parameters — for HPC-specific tasks? More importantly, we believe that such domain-specific LMs would be computationally as well as financially efficient to train.
In this paper, we hypothesize that HPC-specific LMs (e.g., smaller LMs that are designed and trained specifically on HPC datasets) would perform better than existing LMs for HPC tasks. We perform two experiments to validate this hypothesis. The purpose of the first experiment is to build a smaller LM that performs equally well as existing LLMs on generic perplexity score. Towards that end, in the first experiment, we build a smaller LM, called MonoCoder, by reducing the number of layers of PolyCoder code LM by a factor of 4 and pretraining on merely C and C++ codes. We then empirically validate that MonoCoder, despite being a smaller model than PolyCoder, achieves a comparable perplexity score to PolyCoder on the HPCorpus dataset (HPC-specific dataset of C and C++ programs obtained from GitHub). The purpose of the second experiment is then to evaluate the performance of MonoCoder on HPC-specific tasks. Towards that end, in the second experiment, we begin with CodeBLEU competence for high-performance and parallel code generations given increasing context. Subsequently, we finetune MonoCoder and other existing LMs on the dataset of OpenMP parallelization task – pairs of for loops and corresponding OpenMP pragmas. We then evaluate the ability of these models to predict OpenMP private and reduction clauses accurately, including their corresponding variables. We found that existing code LMs capture “local” semantics of code structures (e.g., variable named i is an index variable of a for loop), which leads to their degraded performance on this test. We address this limitation in MonoCoder by developing a novel code pre-processing scheme called Tokompiler, which eliminates any local semantics that the LLM may capture. We found that MonoCoder outperforms existing code LMs in virtually all those tests and tasks, with and without applying Tokompiler. Moreover, and in contrast to MonoCoder, the other LMs’ performances considerably degraded when Tokompiler was applied, suggesting their limited understanding under semantic-less and uncommon settings.
Contributions. We make the following contributions in the paper:
  • By drawing insights from the limitations of existing LLMs, we propose a domain-specific model called MonoCoder for solving tasks related to high-performance computing.
  • We suggest a pre-processing method, called Tokompiler, for evaluating models’ code structure understanding by eliminating local semantics.
  • We compare the proposed approach against PolyCoder and GPT-3.5 — state-of-the-art LLMs — for tasks related to HPC programming.
  • We evaluate MonoCoder on (1) CodeBLEU competence for high-performance and parallel code generations given increasing context and the downstream task of (2) automatic shared memory parallelization via OpenMP, that highlights the strength of the domain-specific model in solving HPC-related tasks.
The remainder of the paper is organized as follows: In section II, we provide motivating examples that underscore the need to build domain-specific LMs. In section III, we unfold the key aspects of HPC programming and how they differ from general programming. Following this, section IV delves into the details of the HPC code dataset HPCorpus and preprocessing steps that should maximize code understanding from the LM. Then section V provides an in-depth exploration of MonoCoder, our language model pre-trained on HPCorpus. section VI introduces Tokompiler, our novel code pre-processor designed for HPC code processing in LMs, detailing its unique approach to generating anonymizing code snippets based on Abstract Syntax Trees (AST). Finally, in section VII, we perform evaluations of MonoCoder and state-of-the-art models on non-HPC and HPC tasks, while in section VIII, we present our efforts in fine-tuning MonoCoder for a specific HPC task, OpenMP pragma generation, demonstrating the practical implications of our work. We discuss related work in section IX. Finally, we conclude the paper in section X and outline directions for future research in section XI.
II Motivations for Building Domain-specific LMs
Recent work has proposed that domain-specific LMs trump general-purpose ones for many use cases. For example, [30] recently showed that a smaller LM trained solely on StackOverflow data outperforms larger models on comprehension tasks such as answer quality prediction and obsolete post detection. In non-software domains, others have shown that domain-specific pretraining can benefit comprehension tasks in biomedicine [31] and finance [32]. It stands to reason that domain-specific pre-training of code LMs may benefit HPC tasks. Here we discuss two specific examples that highlight the limitations of existing general-purpose code LMs.
(a) Correct answer by ChatGPT
(b) Incorrect answer by ChatGPT
Figure 1: ChatGPT example of OpenMP Parallelization Task: “Can this code be parallelized”.
Evaluating ability of GPT-3.5 on an HPC task. Given the popularity of ChatGPT and its ability to assist in programming tasks, we decided to experiment with the underlying GPT-3.5 model [^1] We could not use GPT-4 — an improved version of GPT-3.5 — as it was a premium- and paid-only version. via ChatGPT on the task of parallelizing simple for loops written in C language [15, 26]. We found that (0(a)), ChatGPT could correctly determine the parallelization potential of simple loops (with trivial access patterns) and also provide correct explanations of its determinations [^2] Because GPT-3.5 could parallelize simpler loops with correct explanations, it is reasonable to conclude that its training data included examples of parallel programs written in C language. But when we presented a trickier loop with a non-trivial access pattern (as shown in 0(b)), ChatGPT parallelized the loop incorrectly while also presenting an explanation that was contrary to the prior explanation.
Evaluating ability of StarCoder on a code completion task. Given that GPT-3.5 was designed to solve various tasks, in this experiment, we decided to use an LLM that we specifically developed and trained for programming tasks. Towards that end, we decided to use StarCoder [33], which has caught significant attention in the programming community recently. StarCoder models are 15.5B parameter models pretrained on 80+ programming languages from GitHub (1T training tokens).
Figure 2: An example showing code completion output of StarCoder. Black text shows the input prompt, blue is the generated completion.
In this experiment (Figure 2), we asked the StarCoder model via the HuggingFace website API [^3] https://huggingface.co/bigcode/starcoder to complete a simple C code by providing int main (int argc, const char *argv[]) as the starting prompt. We expected the model to produce an example of “Hello World” written in C or C++. However, to our surprise, it produced strange C code, followed by Python code! We can give a point to the model for producing C code (even though weird-looking), but it is difficult to comprehend why the model produced Python code. Further exploration revealed that the StarCoder model is finetuned on Python from the pretrained, multilingual StarCoderBase model. Naturally, StarCoder has a strong inclination towards Python, which explains why it produced Python code when given C code. Nevertheless, the creators claim that StarCoder “maintains strong performance on other programming languages.” [^4] https://huggingface.co/spaces/bigcode/bigcode-playground. Overall, we believe this example demonstrates that pretrained, multi-lingual models fine-tuned to other datasets can still get confused about the intended language of the input code.
These examples underscore the need for domain-specific training on particular tasks (as opposed to training on several tasks) with a limited number of input languages. This is precisely what we aimed to achieve in this exploration.
III Characteristics of General vs. HPC Programmers
The roles of general programmers and HPC programmers diverge significantly in terms of their tasks, optimization priorities and techniques, tool sets, and development environments [28, 12]. These distinctions arise from the unique demands of HPC programming, where computational speed and efficiency are paramount. In this section, we briefly talk about specific differences that are important in the context of building LLMs.
  1. 1.
    Nature of Tasks: General programmers typically engage in a diverse array of software development tasks, ranging from web and mobile application development to database management and scripting. In contrast, HPC programmers specialize in addressing highly complex scientific and computational problems. Their domain of expertise encompasses tasks such as weather forecasting [34], fluid dynamics simulations [35], structural analysis, and molecular modeling. HPC programmers are deeply involved in the development of applications that demand intensive mathematical modeling and parallel processing capabilities.
  2. 2.
    Performance Optimization: General programmers typically prioritize aspects like code readability, maintainability, and the expeditious development of software. While they acknowledge the importance of performance, it often takes a secondary role in their development efforts. Conversely, HPC programmers place performance at the forefront of their concerns. They meticulously optimize their code to achieve the utmost computational speed. This optimization entails parallelizing code, fine-tuning algorithms, and tailoring code for hardware architectures.
  3. 3.
    Tools and Libraries: General programmers work with a wide spectrum of programming languages and libraries, selecting them based on project requirements. Common programming languages in this context include Python, JavaScript, Java, and C#. In contrast, HPC programmers employ specialized tools and libraries, uniquely suited for scientific and parallel computing tasks. These encompass Message Passing Interface (MPI) [36] for distributed computing, OpenMP [37] for shared-memory parallelism, Intel’s HPC toolkit [38], and CUDA for GPU programming [39]. Furthermore, HPC programmers often work with domain-specific libraries tailored for scientific simulations, with C and C++ being the dominant programming languages.
  4. 4.
    Parallelism and Concurrency: General programming often involves the creation of single-threaded or multi-threaded applications, with parallelism not always being the central concern. Developers may employ threading or asynchronous programming to manage concurrency. HPC programming, conversely, is fundamentally rooted in parallelism. HPC programmers design their applications from the ground up to efficiently leverage parallel processing capabilities [38]. They commonly collaborate with clusters [40], supercomputers, and GPUs to harness maximum computational power.
  5. 5.
    Performance Profiling and Optimization: General programmers employ profiling tools to identify performance bottlenecks in their code. Their optimization efforts may include algorithmic improvements, code refactoring, or selecting appropriate libraries. HPC programmers, on the other hand, rely on advanced profiling and optimization tools such as Intel VTune [41] or NVIDIA Nsight [42] to scrutinize code performance at a granular level. They may engage in comprehensive code rewrites to minimize communication overhead, maximize vectorization, and optimize memory access patterns.
IV HPCorpus: HPC Code Corpus
[…]
Repos Size (GB) Files (#) Functions (#)
C 144,522 46.23 4,552,736 87,817,591
C++ 150,481 26.16 4,735,196 68,233,984
[…]
V MonoCoder: An HPC-specific code LM
MonoCoder. To create a domain-specific model for HPC, we pre-trained a decoder-only transformer model on the language modeling objective (i.e., given past tokens as context, predict the next token in the code) using only C and C++ code from HPCorpus. We excluded the Fortran portion of the dataset given its relatively small size (Table I) and its overall dissimilarity to C/C++. We named this domain-specific model MonoCoder, in contrast to the multilingual Polycoder.
Model size. In the pursuit of deploying domain-specific small language models on average computer systems, careful consideration of model size becomes paramount. To ensure compatibility with such resource constraints, we aimed to design a model that strikes a balance between complexity and memory efficiency. The choice of a model comprising just under 1B parameters emerged as an optimal compromise. The decision is underpinned by the following calculations: Assuming each parameter is represented as a 32-bit floating-point number requiring 4 bytes, the formula RAM size (in bytes) = Number of Parameters $\times$× Size of one parameter (in bytes). Consequently, for 0.9B parameters, the resulting RAM size amounted to 3.6GB. This judiciously selected model size not only accommodates a 4GB RAM constraint – the maximum memory per core in HPC systems [45] – but also leaves additional headroom for other computational processes within the system to accommodate model inference. Note that this worst-case analysis did not consider inference-time optimizations such as quantization, pruning, etc., which would reduce memory consumption further.
We constructed MonoCoder model of 0.9B parameters by reusing the PolyCoder 2.7B model architecture  [27], but reducing the number of layers. This led to an 8-layer model with a hidden dimension of 2560 and 32 attention heads per layer. We used the same tokenizer vocabulary (50,257 tokens). This resulted in an 889.3M parameter model ($\approx$ 0.9B parameters).
Limitations. Our domain-specific model inherits the principles of a Left-to-Right Language Model, which is especially amenable to code generation tasks. By adhering to the left-to-right nature, our model aligns with the established models such as CodeGPT (124M) [46], CodeParrot (1.5B) [47], GPT-Neo [48] and PolyCoder (2.7B) [27], GPT-J (6B) [49], Codex (12B) [50], StarCoder (15.5B) [33], and GPT-NeoX (20B) [51], known for their proficiency in code completion.
However, we acknowledge the inherent challenge posed by the left-to-right approach, which limits the model’s ability to consider context beyond the immediate token sequence. As part of our ongoing research, we are exploring strategies to address this limitation and further refine MonoCoder’s capability to leverage broader contextual information, thereby advancing its effectiveness in the nuanced landscape of HPC code synthesis. This will be demonstrated especially in the design of our downstream tasks in section VIII.
Figure 3: MonoCoder training and validation loss curves.
Pre-training details. We pre-trained the MonoCoder model using the GPTNeoX framework [52] on 4 NVIDIA A40 48GB GPUs with fp16 precision. For pre-training, we used the Adam optimizer and followed a learning rate schedule with a linear warmup for the first 1% of steps and a cosine decay over the remaining steps. Gradients were clipped at 1.0. Each training sample had a maximum length of 2048 tokens and was trained in mini-batches of 16 samples (4 per GPU). The model was trained for 160K steps at a learning rate of 0.00008. Figure 3 shows the training and validation loss curves of MonoCoder.
VI Tokompiler: Preprocessing to Evaluate Code Structure Understanding
(a) GPT3 Tokenizer
(b) Codex Tokenizer
(c) GPT-3.5 & 4
(d) Tokompiler
Figure 4: GPT3, Codex, GPT-3.5&4 and Tokompiler tokenization of the Pi C code. As models are enhanced, tokenization is also improved, but it is still imperfect for domain-specific language primitives and numbers.
// Source code:
int main() {
  int r[2800 + 1];
// Tokompiler:
int func_252() {
 int arr_88[num_34 +
            num_842];
// Lexicalized tokens:
["int", "func", "_", "252",
"()", "{", "int", "arr", "_"
... (tokens continue)
Figure 5: Tokompiler pipeline overview: Given a source code, the code turns into a semantic-less version using AST knowledge, and eventually, the lexicalized tokens are fed into MonoCoder.
As detailed in section III-section IV, it is crucial that an effective code LM for HPC tasks understands both the syntax and structure of the code, without relying on the memorization of potentially misleading human semantics [53] (e.g., variable names). We propose preprocessing code based on parsing the abstract syntax tree (AST) to remove semantic information. A code LM that performs well on this anonymized preprocessed code is clearly relying on its understanding of code structure, rather than natural language-derived semantics. We name this preprocessing method Tokompiler, inspired by input tokenization and by ensuring functionally correct and compilable code.
Tokenizing code for LLMs necessitates specialized techniques to accommodate programming language syntax. [^5] These approaches include utilizing BPE and subword tokenization akin to natural language [54], employing syntax-aware tokenization to identify language-specific elements like keywords and identifiers [55], constructing tokens based on the AST to capture structural information [56], implementing language-specific lexers following grammar rules [57], preserving character integrity with character-level tokenization [58], tailoring tokenization to unique syntax rules, and leveraging dedicated code tokenization libraries. LLMs geared towards code comprehension, such as GPT-3.5-Turbo for code, likely combine several techniques, prioritizing syntax-aware tokenization to effectively process and generate code snippets in various programming languages and tasks [59]. An examination of tokenization strategies in GPT models reveals some deficits. Figure 4 presents an example of GPT3 (3(a)), Codex (3(b)), and GPT-3.5&4 (3(c)) tokenization of the Pi C code. In GPT3, one of the issues is that each individual whitespace is a new token. An advanced pre-processor is used for Codex and GPT-3.5&4. However, it is still unaware of the need to keep language primitives and numbers intact; instead, it splits numbers into subtokens.
Tokompiler preprocessing (Figure 5) replaces variable names, numbers, and strings with random variable names and removes superfluous input (e.g., comments, extra whitespace). In detail:
  1. 1.
    Generate Replaced Code: Create a version of the original code with anonymized variable names, numbers, and strings. The intuition behind this step is to eliminate misleading semantics, such as variable i being an index variable of a for loop in the C language.
  2. 2.
    AST Generation: Parse the anonymized code using TreeSitter [^6] https://github.com/tree-sitter/tree-sitter or any suitable parser to generate an AST.
  3. 3.
    AST to Code: Transform the updated AST back into code, thus eliminating any comments that may interfere with anonymization. Although the removal of natural language (NL) from code may hamper the model’s ability to solve NL tasks, compilation tasks do not require NL understanding.
  4. 4.
    Random Number Attachment: For recurrent tokens (e.g., var_1 or num_2), use statistics to attach random integers from a predefined range (e.g., 1 to 1000) during tokenization. The attached numbers are randomly chosen without any relation to the type or order of the replaced tokens or the file/function length. This step also eliminates misleading semantics. For instance, if variable i is consistently replaced with var_1, then the model may learn that var_1 is an index variable of for loops.
VII MonoCoder Evaluation
VII-A Language Modeling Evaluation
{tikzpicture}
(a) Model Size
{tikzpicture}
(b) Perplexity C
{tikzpicture}
(c) Perplexity C++
Figure 6: Perplexity comparison across code LMs. MonoCoder was evaluated on the code from the test set of HPCorpus. The data from the other models are from [27] and [33].
We start by measuring the perplexity of MonoCoder and other LLMs on unseen test programs in C and C++.
Metric: Perplexity. Perplexity is a metric commonly employed in natural language processing and language modeling to assess the efficacy of a probabilistic model in predicting a given sequence of tokens. It serves as a measure of the model’s uncertainty or confusion when assigning probabilities to a set of observed data. A lower perplexity value signifies the better predictive performance of a model. In the context of our research, we measured perplexity to evaluate the predictive capabilities of the pre-trained MonoCoder language model. That is, this assessed how well the model assigned probabilities to the token sequences in the C and C++ portions of the HPCorpus test set. We then compared this perplexity to various pre-trained LLMs assessed on other C and C++ datasets in published work [27, 33]. The StarCoder values are for the model with a 2K context window, similar to MonoCoder’s 2048 context window.
Setup. For this test, we use the training procedure outlined earlier to pre-train MonoCoder on HPCorpus. We then measure perplexity on C and C++ code from the test set of HPCorpus. We compare these values to published perplexity measurements of other, larger pre-trained code LMs from other C and C++ test data.
Results. The perplexity comparison can be seen in Figure 6. We can see that despite the much smaller size of MonoCoder, it only performs slightly worse than much larger models. On C language, it is performing better than the 3x-larger GPT-Neo model.
VII-B Code Generation
In this second evaluation, we assess the ability of MonoCoder and other LLMs in generating the remainder of a code block when provided with varying amounts of prior context. While perplexity provides information about the uncertainty of the model, it does not necessarily measure the quality of the generated code. For this reason, we further evaluate code understanding with the CodeBLEU [60] score.
Metric: CodeBLEU score. In the realm of code synthesis, the selection of an appropriate evaluation metric plays a pivotal role in discerning the efficacy of machine-generated code. Commonly utilized metrics, such as BLEU [61], while effective in natural language contexts, fall short in comprehensively evaluating code quality due to their inability to account for syntactic and semantic intricacies inherent in programming languages. To address these limitations, we have chosen to employ the CodeBLEU [60] metric in our evaluation framework. CodeBLEU amalgamates the robustness of BLEU, incorporating n-gram matching, with an innovative integration of code syntax and semantics through AST and data-flow structures. This holistic approach provides a nuanced evaluation that extends beyond token matching, considering the significance of keywords, syntactic accuracy, and semantic correctness. In the case of HPC code without natural language, CodeBLEU is the appropriate metric to check similarity to generations.
Setup. For this evaluation, we first devise two sub-datasets: (1) General dataset of about 20K examples of C and C++ programs from HPCorpus with HPC-orientation (ones that used to contain parallel for pragmas. However, those pragmas were intentionally removed to keep clean C and C++ programs that have HPC orientation) and (2) OpenMP dataset of about 20K examples of C and C++ programs containing OpenMP code (e.g., parallel for). The method to test the model’s understanding is by incrementally supplying it parts of those programs (100, 300, and 600 first tokens, out of an average of 1200 tokens per code) and evaluating the outcome of the rest of the function using CodeBLEU. Demonstration of the idea is presented in Figure 7.
1int main (int argc, char *argv[]) {
2// COMPLETION:
3int i;
4int n = argc - 1;
5int m = argc - 2;
6int k = argc - 3;
7int l = argc - 4;
8int r = argc - 5;
9int c = argc - 6;
10int d = argc - 7;
11int e = argc - 8;
12int f = argc - 9;
13int g = argc - 10;
14int h = argc - 11;
15int i1 = argc - 12;
16int i2 = argc - 13;
17int i3 = argc - 14;
(a) Given function signature
1int main (int argc, char *argv[]) {
2int i,j;
3double x;
4double pi, sum = 0.0;
5double start, delta;
6step = 1.0/(double) steps;
7for (j=1; j<= MAX_THREADS; j++) {
8printf("running on %d threads: ", j);
9omp_set_num_threads(j);
10// COMPLETION:
11start = omp_get_wtime();
12for (i=0; i<steps; i++) {
13x = (double) i;
14pi = 3.141592653589793238
154626433832795028841971693
169937510582097494459230781
176406286208998628034825342...
(b) Given initialization
1int main (int argc, char *argv[]) {
2int i,j;
3double x;
4double pi, sum = 0.0;
5double start, delta;
6step = 1.0/(double) steps;
7for (j=1; j<= MAX_THREADS; j++) {
8printf("running on %d threads: ", j);
9omp_set_num_threads(j);
10sum = 0.0;
11double start = omp_get_wtime();
12#pragma omp parallel for reduction(+:sum) private(x)
13// COMPLETION:
14for (i=0; i<steps; i++) {
15    x = (double) i;
16    sum += x;
17} ...
(c) Given context of parallel execution
Figure 7: Testing generation of HPC code by the foundation models (PolyCoder example): Evaluating machine-generated text given different contexts of the initial HPC codes and measures the similarity to the true reference.
Results. 7(a) shows code completion performance of the models on General dataset. For a context of the first 100 tokens, MonoCoder achieves a CodeBLEU score of 0.62, while PolyCoder and GPT-3.5 attain scores of 0.45 and 0.44, respectively. This discrepancy can be attributed to PolyCoder’s diverse training across programming languages and GPT-3.5’s general nature, resulting in versatile but less HPC-specific answers (the limitations highlighted in section II). Expanding the context window to 600 tokens further widens the CodeBLEU score gap between MonoCoder and GPT-3.5, emphasizing lack of HPC knowledge in GPT. Overall, the model’s reduced orientation towards HPC codes accentuates the discernible gap in HPC code knowledge among the models.
{tikzpicture}
(a) Code completion performance on General dataset.
{tikzpicture}
(b) Code completion performance on OpenMP dataset.
Figure 8: Performance of various models on Code Completion task.
Similar trends are evident on OpenMP dataset in 7(b), depicting the CodeBLEU scores of models when completing functions containing OpenMP pragmas. Generally, the results are slightly lower than the those on General dataset, underscoring the models’ limited understanding of OpenMP.
It is noteworthy that, despite the Tokompiler pre-processing having a minor impact on the performance of MonoCoder and GPT-3.5, this code representation significantly impairs the performance of PolyCoder. This observation suggests that PolyCoder heavily relies on local semantics when completing programs.
VIII Downstream HPC Task: OpenMP Pragma Generation
In this section, we evaluate the ability of MonoCoder and other existing LLMs in assisting on HPC specific downstream task. Specifically, we evaluate ability of these models in automatically suggesting appropriate OpenMP pragma for the input C/C++ program.
Task objective. While a lot of progress has been made in integrating machine learning into the identification of parallelism opportunities in C/C++ programs [15, 26, 62], there has been a notable gap in connecting parallelism detection with code generation. In an effort to bridge this gap, our work leverages MonoCoder, refining its capabilities through fine-tuning on the OpenMP-specific subcorpus of HPCorpus. Specifically, we created OMPify+ by fine-tuning MonoCoder to generate complete OpenMP pragmas, when presented with a for loop written in C/C++ languages. OpenMP pragma annotations on for loops allow such loops to improve performance by using multi-threading to process computations inside the loop concurrently.
Dataset. To create a dataset for automatic OpenMP parallelization task, we curated a subcorpus of HPCorpus named HPCorpusOMP . This subcorpus specifically includes all the for loops that have either #omp parallel for pragma that distributes the workload across multiple CPU cores, or offload tasks to a team of threads for GPU execution using #omp target teams distribute. The dataset also includes OpenMP pragmas containing private and reduction clauses. The details of the dataset in terms of these clauses is present in Table II.
private reduction target Total
C 3,526 713 145 8,764
C++ 2,122 1,424 345 18,151
TABLE II: Clause breakdown in HPCorpusOMP .
It is important to note that we excluded context of the loops to maintain focus on the loop-specific information. However, to avoid the loss of crucial details, we designed HPCorpusOMP such that the private clause encompasses instances of both firstprivate and lastprivate, as both of these clauses serve the purpose of creating a private instance of variables for each thread.
Evaluation setup. We finetuned MonoCoder on the HPCorpusOMP subcorpus. The finetuning task was setup as a generation task in which the complete for loop is input to the model, and the expected generated output was the pragma and its associated variables. We then evaluated MonoCoder on a test set of HPCorpusOMP , both with the original for loops and their anonymized versions using Tokompiler. For comparison, we assessed GPT-3.5 turbo in a zero-shot manner. Specifically, we supplied a prompt instructing it to “Generate the optimal OpenMP pragma for the provided code” when presented with the for loop.
Fine-tuning details. We fine-tuned the pretrained MonoCoder model to the OpenMP pragma generation task using the Huggingface transformers library on 2 NVIDIA V100 32GB GPUs at fp32 precision. Fine-tuning used the AdamW optimizer with a linear warmup over the first 100 steps, followed by a linear decay over the remaining steps. Fine-tuning samples also had a maximum length of 2048 tokens, trained in 16 sample minibatches (8 per GPU). The fine-tuning was run for 4 epochs at the learning rate of 0.000016.
(a) Clause evaluation.
(b) Inner-clause evaluation.
Figure 9: Classification tasks for pragma evaluation.
{tikzpicture}
(a) Performance on classifying private clause (Task \@slowromancapi@).
{tikzpicture}
(b) Performance on classifying reduction clause (Task \@slowromancapii@).
Figure 10: Performance of MonoCoder and GPT on OpenMP pragma classification tasks.
Figure 11: Generation of private clause (Task \@slowromancapiii@).
{tikzpicture} (a) Accuracy of predicting full private clause. {tikzpicture} (b) Cumulative distribution of samples.
{tikzpicture} (a) Accuracy of predicting full reduction clause. {tikzpicture} (b) Cumulative distribution of samples.
Figure 11: Generation of private clause (Task \@slowromancapiii@).
Figure 12: Generation of reduction clause (Task \@slowromancapiv@).
Evaluation method. While we used the CodeBLEU score to evaluate the fluency of generated code containing OpenMP pragmas, it falls short in capturing correctness aspects. To comprehensively evaluate the performance of MonoCoder and GPT, we broke down the performance measurement of the OpenMP pragma generation task into five subtasks:
  1. 1.
    Clause-level generation: This set of tasks evaluates whether the model correctly generates OpenMP private (Task \@slowromancapi@) and reduction (Task \@slowromancapii@) clauses. For example, in the sample depicted in 8(a), the evaluation involves marking true-positive (TP) for the private clause as it is present in both the label and prediction. Conversely, for the reduction clause, we mark it as false-negative (FN) as the model fails to generate the required clause present in the label. We also report precision and recall to account for true negatives (TN), since many samples do not contain private or reduction clauses. For the reduction operator, we will focus solely on accuracy, since the result is deemed correct if the operator is the same and incorrect otherwise. The results below depict the outcomes of these evaluations.
  2. 2.
    Generating the correct variables of OpenMP clauses: This set of tasks evaluates the correctness of the generated variables for each private (Task \@slowromancapiii@) and reduction clause (Task \@slowromancapiv@). For the reduction clause, we also measure the accuracy of generating the reduction operator (Task \@slowromancapv@). Our evaluation recognizes that clause variables, like clauses themselves, could be unordered. 8(b) illustrates the evaluation method with an example. For example, variables A and B would be considered true-positive (TP), while variable X would be a false-positive (FP), as it is present in the generated code but not in the ground-truth label.
Results on clause-level generation. In Task \@slowromancapi@, MonoCoder provided accuracy of 94.4%, surpassing GPT 3.5, which achieved 77% (9(a)). The substantial gap in recall between MonoCoder and GPT can attributed to GPT frequently mispredicting the need for the private clause (FN), thereby impacting its overall performance. Similarly, in Task \@slowromancapii@, MonoCoder exhibited an accuracy of 98.3%, outperforming GPT with an accuracy of 92.3% (9(b)).
An intriguing observation is the minimal impact of the anonymized Tokompiler version on MonoCoder’s results, indicating robustness. In contrast, GPT’s performance significantly degraded with the anonymized version, suggesting its reliance on local semantics rather than the underlying code structure, unlike MonoCoder. These findings emphasize MonoCoder’s ability to leverage code structure for accurate predictions in contrast to GPT, which heavily depends on local semantic information.
Results on generating clause variables. OpenMP private and reduction clauses can contain different number of variables. To understand the impact of different number of variables, we report performance of Task \@slowromancapiii@ and \@slowromancapiv@ in terms of the number of variables in those clauses.
In Task \@slowromancapiii@ (Figure 12), the analysis reveals that the majority of private clauses involve up to five variables. For private clauses with one variable, MonoCoder generates 69% of the variables, while GPT succeeds only with 5.3%. When examining all the clauses, regardless of their variable count, MonoCoder accurately generates 48.5% of these variables, while GPT scores only 14.3%. We can make two key observations from these results. Firstly, for codes that do not include many potential variables that should be private, it is easier to identify the variables that should be private to each thread. Furthermore, the substantial performance gap between GPT and MonoCoder in this examination underscores a significant issue: GPT often fails to generate variables within the clause, even when they are present in the labeled clause.
Similar trends are observed in Task \@slowromancapiv@ (Figure 12). For reduction clauses with a single variable, MonoCoder achieves 58.1% accuracy, while GPT achieves 55.2%. When examining all the reduction clauses, MonoCoder attains an accuracy of 51.7%, while GPT achieves 51.2%. Notably, the anonymized Tokompiler version has a minimal impact on MonoCoder’s variable generation. In stark contrast, the anonymized version significantly degrades GPT’s performance, underscoring its deficiency in comprehending code structure.
In Task \@slowromancapv@, MonoCoder generates the reduction operator correctly 81% of the time, while GPT succeeds 79% of the time. The similar performance between GPT and MonoCoder, along with the minimal impact of the anonymized Tokompiler version on both the models, can be explained by the inherent simplicity of the task. The detection of the reduction operator becomes relatively straightforward when there is a variable aggregating the computation along the loop iterations.
IX Related Work
The code LM literature is somewhat divided into NLP-oriented approaches and software engineering approaches. The NLP-oriented approaches typically evaluate code LMs with perplexity and datasets like HumanEval [50] or Mostly Basic Programming Problems [63], which assess both natural language comprehension and general reasoning ability along with code comprehension. Evaluations more specific to code include tasks such as code search or code repair [64]. Only very recently have we seen an evaluation of code LMs for HPC-specific tasks, such as OpenMP pragma prediction [16] and parallel code generation [29], and race detection [28]. Our work complements these recent papers by demonstrating how a smaller, domain-specific LM can accomplish HPC tasks. We also demonstrate a novel way to assess how heavily a model relies on local semantics, rather than a structural understanding of the code, to perform various tasks such as code completion.
While we have mentioned work on domain-specific LMs [30, 31, 32] in section II, there are also related investigations in NLP on improving the reasoning and few-shot learning capabilities of smaller LMs [65, 66, 67]. These investigations designed procedures tailored to the task of interest to enable smaller models to perform as well, if not better than, larger ones. Our work contributes to a general trend in improving LMs for use in low-resource settings.
X Conclusion
Large language models, such as GPT-3.5, GPT-4, etc., have revolutionized the field of natural language processing, with their chat-based versions, such as ChatGPT, catching significant attention worldwide. This has naturally led to their applications in programming and, specifically, high-performance computing. Our experiments with existing LLMs, however, revealed that these LLMs cannot solve HPC-related programming problems accurately, or they are multi-lingual and are trained on languages that are not HPC-related. More importantly, these models are huge in size and demand expensive compute resources to train.
That is why, in this paper, we decided to evaluate if we can build smaller models that can do similar, if not better, than existing LLMs on HPC-related programming tasks. We built MonoCoder using existing code-oriented LLM but by carefully selecting model size such that the resulting model can fit on commodity hardware on edge devices. Our experimental results demonstrate that such smaller models can be trained on HPC-specific languages and can perform similarly to existing LLMs in terms of perplexity and code completion tasks. Furthermore, after fine-tuning MonoCoder on the dataset containing OpenMP pragmas, we found that it can outperform the considerably larger GPT-3.5 model on the task of generating OpenMP pragmas. We hope that our results will help in democratizing applications of smaller LLMs to HPC programming problems.
XI Future Work
In the near future, we intend to apply Tokompiler to C and C++ corpora and integrate more code representations, such as the data-flow graph (DFG) and the intermediate representation (IR) [68], to enhance model understanding as shown in closely related works  [69, 70]. Moreover, we believe that pre-training on the provably compilable subset of HPCorpus will enhance the correctness of model predictions for compilation-oriented tasks. Then, we intend to fine-tune those pre-trained models for HPC downstream tasks, such as OpenMP pragma generation [16, 15] and MPI domain decomposition distribution [16, 17]. In general, given the differences between general programming tasks and HPC-specific programming tasks, our research vision is to systematically analyze each and every element of existing LLMs (model architecture, dataset, etc.) and redesign them as needed for HPC-specific tasks.